OpenStudy (anonymous):

Hey. This is my first time using OpenStudy. Here's my two questions (from Calc III): 1. Let P = (1, 2, -1). Find the point of intersection of the plane 3x-4y+z=2, with the line through P, perpendicular to that plane. 2. Let Q = <1, -1, 2>, P = (1, 3, -2), and N = <1, 2, 2>. Find the point of the intersection of the line through P in the direction of N, and the plane through Q perpendicular to N. Can you show me how to work the problems out? Thanks!

OpenStudy (anonymous):

Does anyone have an idea of how to approach this question?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

pull the coeffs off your plane eq for me ... what are they?

OpenStudy (anonymous):

No plane coeffs are given, what you see is what the question asks. N = the Normal. P = a Point, Q = a vector.

OpenStudy (amistre64):

maybe this might be more productive; set up a vector using the coeffs and variables from the plane eq as components

OpenStudy (amistre64):

3x-4y+z=2 <3x,-4y,1z> now take away the variabels <3,-4,1> this is the vector that is perp to the plane; this is the normal to the plane

OpenStudy (amistre64):

use this vector and anchor it to the point given so that we can define all x y and zs on the line

OpenStudy (anonymous):

So would I set it up like this: x = 1 + 3t y = 2 -4y z = -1 +1z

OpenStudy (anonymous):

***x = 1 + 3t y = 2 -4t z = -1 +1t

OpenStudy (amistre64):

yes, very good :) now we know what x y and z has to equal for any given point; and we want to know which point is in the plane itslef

OpenStudy (amistre64):

substitute you xyz eqs into the variables of the plane and solve for t

OpenStudy (anonymous):

What variables am I substituting for t?

OpenStudy (anonymous):

or x, y, z rather

OpenStudy (amistre64):

knowing: x = 1 + 3t y = 2 -4t z = -1 +1t then the point in the line in the plane is satisfies: 3x -4y +z = 2 3(1+3t) - 4(2 -4t) + (-1 +1t) = 2 solve for t to know how to find the xyz point

OpenStudy (anonymous):

oh awesome, thanks! - simple enough. how about the next question?

OpenStudy (amistre64):

lets see Q = <1, -1, 2> P = (1, 3, -2) N = <1, 2, 2> Find the point of the intersection of the line through P in the direction of N, and the plane through Q perpendicular to N. I take it P is a point and QN are vectors

OpenStudy (amistre64):

set up the line then using P as the anchor and N as the direction vector

OpenStudy (amistre64):

x= Px +Nx t y= Py +Ny t z= Pz +Nz t

OpenStudy (amistre64):

the plane thru Q?

OpenStudy (amistre64):

perp to N means N is our normal vector; do we include Q in the plane? hard to see what its asking for clearly

OpenStudy (anonymous):

x = x + t y = 3y + 2t z = -2z +2t

OpenStudy (anonymous):

yeah I understand the question is worded a bit weird. it's from a textbook. if it helps, this is the answer: (1, 3, -2) - I just want to know how that answer was gotten

OpenStudy (amistre64):

 is what im thinking but there is no point to anchor any of this to

OpenStudy (amistre64):



OpenStudy (amistre64):

in fact we can slide this plane up and down the N vector; Im missing a way to read how to get a point to anchor the plane to

OpenStudy (anonymous):

Well, the question wants us to find a point, and the answer is, according to the book, (1, 3, -2). Interesting to note that that is the same as the point given

OpenStudy (amistre64):

(1,3,-2) is on one of an infinite number of planes

OpenStudy (amistre64):

yeah, but there is no way to determine that from my point of view other than saying P is in the plane

OpenStudy (amistre64):

and it doesnt specify ANY point to anchor this thing down with

OpenStudy (anonymous):

yeah what I wrote initially is all it says. i see why it's an odd question though

OpenStudy (amistre64):

read the question again and make sure you havent left anything out

OpenStudy (anonymous):

2. Let Q = (1, -1, 12), P = (1, 3, -2), and N = (1, 2, 2). Find the point of the intersection of the line through P in the direction of N, and the plane through Q perpendicular to N.

OpenStudy (anonymous):

maybe use of the cross product is required?

OpenStudy (amistre64):

ok, those are 3 points in space in that notation; which means we find the line running thru N and P